problem #1
How many kilograms of each of a 50% silver alloy and 10% silver alloy must be mixed to obtain 24 kilograms of a 25% silver alloy?
problem #2
Two vats contain cream, one sample tested 8% butterfat and the other tested 16% butterfat. How much cream should be taken from each to produce a 5 liters mixture which will test 12% butterfat?
problem #3
How many kilograms each of a powdered detergent and a rinsing solution containing 80% of the pure rinsing solution must be mixed to make 140 kilograms of a 24% detergent?
I'll be waiting for the answer and solution

Answer by lwsshak3(11628)

problem #1
How many kilograms of each of a 50% silver alloy and 10% silver alloy must be mixed to obtain 24 kilograms of a 25% silver alloy?
problem #2
Two vats contain cream, one sample tested 8% butterfat and the other tested 16% butterfat. How much cream should be taken from each to produce a 5 liters mixture which will test 12% butterfat?
problem #3
How many kilograms each of a powdered detergent and a rinsing solution containing 80% of the pure rinsing solution must be mixed to make 140 kilograms of a 24% detergent?
...
problem #1
let x=kilograms of 50% silver alloy to be mixed
24-x=kilograms of 10% silver alloy to be mixed
50%x+10%(24-x)=25%(24)
.5x+2.4-.1x=6
.4x=3.6
x=9
24-x=24-9=15
ans:
9 kilograms of 50% silver must be mixed with 15 kilograms of 10% silver to obtain 24 kilograms of a 25% silver alloy?
..
problem #2
let x=liters of 8% butterfat cream
5-x=liters of 16% butterfat cream
8%x+16%(5-x)=12%(5)
.08x+.8-.16x=.6
multiply by 100
8x+80-16x=60
-8x=-20
x=20/8=2.5
5-x=5-2.5=2.5
ans:
2.5 liters of 8% butterfat cream and 2.5 liters of 12% butterfat cream should be taken to produce a 5 liters mixture which will test 12% butterfat
..
problem #3
sorry, I don't understand this problem